OpenStudy (anonymous):

If (log_3 X) (log_5 3) = 3, find x

OpenStudy (anonymous):

Which one of these has an x?

OpenStudy (anonymous):

The other one is just a multiplicative constant. You can divide it over to the other side.

OpenStudy (anonymous):

?? how? the bases are diff-- three and five...how can u just divide?

OpenStudy (anonymous):

Because the log base 5 of 3 is just a number. Like 4/5 or the square root of 15

OpenStudy (anonymous):

=( i still don't get it....

OpenStudy (anonymous):

Ok here. Lets say that \[C=log_5 3\] Now your equation looks like: \[C(log_3 x) = 3\] Can you solve it now?

OpenStudy (anonymous):

hmm..you know how log_b Y = x is equal the y=b^x in exponential form? how could you do that with this equation? isn't that how ur supposed to solve these things?

OpenStudy (anonymous):

Yes, but first you have to get \(log_3x\) by itself.

OpenStudy (anonymous):

so it's log_3 x = 3c and then 3^3c=x

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

whaat???nooo

OpenStudy (anonymous):

C(log_3 x) = 3 So you have to divide by C, not multiply

OpenStudy (anonymous):

OH. lol...right.

OpenStudy (anonymous):

when i solve for (log_5 3)--to conv to exp form, does it still equal to three, which is the what the whole equation is equal to?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

That part is trickier

OpenStudy (anonymous):

But lets see what you have with C for now.

OpenStudy (anonymous):

x=3^3/c

OpenStudy (anonymous):

Is that (3^3)/c or 3^(3/c) ?

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

Correct. Ok, so now to find what C is. \[C = log_5 3 \iff 5^C = 3\]

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ok, so that means that if we take the ln of both sides we get \[C(ln\ 5) = (ln\ 3)\]

OpenStudy (anonymous):

polpak i need ur help please

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

So that means \[C = \frac{(ln\ 3)}{(ln\ 5)}\]

OpenStudy (anonymous):

oh~~~

OpenStudy (anonymous):

Which means that \[log_53 = \frac{(ln\ 3)}{(ln\ 5)}\]

OpenStudy (anonymous):

oh, wow! thanks so much!!

OpenStudy (anonymous):

that makes so much more sense now.

OpenStudy (anonymous):

Glad I could help.